OpenStudy (anonymous):

Find an equation of the tangent line to the curve y = sin(7 x) + cos(4 x) at the point (pi/6,y(pi/6) ). Tangent line:

zepdrix (zepdrix):

Hey Carol c: So we're trying to find a line that is tangent to our curve at a specific point. That line will be of the form:\[\Large\rm y=\color{orangered}{m}x+b\]Where m, the slope, is given by the derivative function evaluated at that particular point.

OpenStudy (anonymous):

Yes I know I would plug in pi/6 to the equation but I keep getting a weird answer

zepdrix (zepdrix):

\[\Large\rm y=\color{orangered}{f'\left(\frac{\pi}{6}\right)}x+b\]

zepdrix (zepdrix):

\[\Large\rm f(x)=\sin(7x)+\cos(4x)\]Did you have any trouble taking this derivative?

zepdrix (zepdrix):

(I'm calling it f(x) instead of y, otherwise we'll confuse it with our tangent line y D: )

OpenStudy (anonymous):

It would be 7cos7x-4sin4x correct?

zepdrix (zepdrix):

mmm good good.

OpenStudy (anonymous):

and then i plug in pi/6

OpenStudy (anonymous):

and i get (-1/2)-(2radical3)

zepdrix (zepdrix):

Hmm I'm coming up with:\[\Large\rm f'\left(\frac{\pi}{6}\right)=7\left(-\frac{\sqrt3}{2}\right)-4\left(\frac{\sqrt3}{2}\right)\]

OpenStudy (anonymous):

Maybe I made an algebraic error. But what would my next step be?

zepdrix (zepdrix):

\[\Large\rm \cos\left(\frac{7\pi}{6}\right)=-\frac{\sqrt3}{2}\]

zepdrix (zepdrix):

That value is the slope of our tangent line. Let's simplify it down before plugging it in.

zepdrix (zepdrix):

So we have -7 of those things, and -4 more of them. So we have -11sqrt3/2, something like that, yah?

OpenStudy (anonymous):

yes thats correct

zepdrix (zepdrix):

So here is what our tangent line currently looks like with our new found slope, \[\Large\rm y=\color{orangered}{-\frac{11\sqrt3}{2}}x+b\]

OpenStudy (anonymous):

that makes sense

zepdrix (zepdrix):

To find b, the y-intercept, we need to be able to plug in a point that lies on the line.

zepdrix (zepdrix):

Since the line is `tangent` or `touching` the curve at one specific point, the curve and tangent line will share that coordinate point.

zepdrix (zepdrix):

But notice that they're giving you a little bit of extra work, they only gave you the x coordinate :)

OpenStudy (anonymous):

thats true

OpenStudy (anonymous):

:(

zepdrix (zepdrix):

So we need to find out what the coordinate point is that the two functions share. We need the y coordinate that corresponds to x=pi/6. So now you'll plug pi/6 into the `original function` to get a coordinate pair.

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

i get -1

zepdrix (zepdrix):

mmm me too, ok good good. So the point we can use for our tangent line is:\[\Large\rm \left(\frac{\pi}{6},-1\right)\]

OpenStudy (anonymous):

yay!

zepdrix (zepdrix):

So plug in the stuff c: what do you get for your b? It might be a little messy looking!

OpenStudy (anonymous):

is my b the -1?

zepdrix (zepdrix):

So far we've establisshed this:\[\Large\rm \color{royalblue}{y}=-\frac{11\sqrt3}{2}\color{royalblue}{x}+b\]And we can use our coordinate point \(\Large\rm \left(\color{royalblue}{\dfrac{\pi}{6}},\color{royalblue}{-1}\right)\) to find our b.

OpenStudy (anonymous):

ahhh i see so then...\[-1=\frac{ -11\sqrt{3} }{ 2 }(\pi/6)\]

zepdrix (zepdrix):

With a +b on the end, yes.

OpenStudy (anonymous):

yes and thenn

OpenStudy (anonymous):

im confused what to do with the pi/6

zepdrix (zepdrix):

Multiply your fractions, bottom with bottom, top with top. The you'll need to do some addition to isolate your b.

OpenStudy (anonymous):

\[-b=-66\sqrt{3}*2\pi\]

OpenStudy (anonymous):

and plus one at the end

zepdrix (zepdrix):

Woah, b should have been a fraction. :o Where did the denominator go?

OpenStudy (anonymous):

I cross multiplied? I'm sorry that I'm not getting this as easy as it should be

zepdrix (zepdrix):

When we multiply fractions, we multiply straight across, yes? So this:\[\Large\rm -1=-\frac{11\sqrt3}{2}\cdot\frac{\pi}{6}+b\]Gives us this: \[\Large\rm -1=-\frac{11\sqrt3 \pi}{12}+b\]From here we're NOT allowed to cross multiply since that +b is hanging out over there. If both sides were completely fractions, then yes we could do so.

zepdrix (zepdrix):

You want to add that ugly monstrous thing to the other side.

OpenStudy (anonymous):

that makes much more sense

OpenStudy (anonymous):

Do i have to simplify further after adding it to the other side or is that my final answer?

zepdrix (zepdrix):

\[\Large\rm b=\frac{11\sqrt3~\pi}{12}-1\]You can get a common denominator if you prefer to look at your b as a single term. But this is fine right here. For calculus it's nice to get in the habit of `not oversimplifying things`.

zepdrix (zepdrix):

Just make sure you realize that this is not your final answer, you need to plug this into your tangent line to get your final answer.

OpenStudy (anonymous):

ok so then...

OpenStudy (anonymous):

\[\frac{ -11\sqrt{3} }{ 2 }x+\frac{ 11 \sqrt{3}}{12 }\]

OpenStudy (anonymous):

theres a pi and -1

OpenStudy (anonymous):

idk why it keeps cutting off

zepdrix (zepdrix):

hehe

zepdrix (zepdrix):

yah that sounds right! \c:/ \[\Large\rm y=-\frac{11\sqrt3}{2}x+\frac{11\sqrt3~\pi}{12}-1\]

zepdrix (zepdrix):

Here is graph of the function so you can see what it looks like and confirm that this crazy mess is correct. https://www.desmos.com/calculator/qrj741tptf

OpenStudy (anonymous):

let me plug in to check right now

zepdrix (zepdrix):

It's kind of hard to read since the slope is so steep, but the blue line is tangent to the red curve at pi/6

OpenStudy (anonymous):

thank you sososo much for your help!

OpenStudy (anonymous):

it means a lot

zepdrix (zepdrix):

yay team \c:/

OpenStudy (anonymous):

hehe

OpenStudy (anonymous):

Since youve been such a good help can you help me with one more problem?